A body is in straight line motion with an acceleration given by a=32-4v . The initial conditions are at t=0 , v=4 . Find the velocity when t=ln2.

Grade:12

1 Answers

mahendra shah
10 years ago

dv/(32-4v)  = dt..

integrate

lhs from 4 to v

rhs from 0 to ln2

thus

-4ln(32-4v)]v,4 = ln2

32-4v]v,4 = 2^(-1/4)

32-4v -32 + 16 = 0.85

4v = 15approx

v = 3.7